OpenStudy (anonymous):

What is the 8th term of the geometric sequence -1, 3, -9, ...?

OpenStudy (cwrw238):

first find common ratio this = -3 /-1 the 8th term is ar^(n-1) where a = first term (-1), r = common ratio and n = 8

OpenStudy (cwrw238):

* sorry common ratio = 3 /-1

OpenStudy (anonymous):

\[\large t_n = a.r^{n-1}\] Here put r = 8, a = -1 and r = -3

OpenStudy (anonymous):

Sorry put n = 8 a = -1 and r = -3

OpenStudy (anonymous):

2187

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Yes you are right.. But there is a little mistake I think..

OpenStudy (cwrw238):

= -1 * (-3)^7 = 2187

OpenStudy (cwrw238):

no mistake

OpenStudy (anonymous):

I was just kidding.. There is no mistake...

OpenStudy (anonymous):

lol mk byeeeeeeeeeeeeee ty